The solution of the equation log 7 log 5 (√x + 5 + √x) = 0 is………….

  1. The solution of the equation log7 log5 (√x + 5 + √x) = 0 is………….

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Hello Student,
Please find the answer to your question

The given equation is
log7 log5 (√x + 5 + √x) = 0
\Rightarrow log5 (√x + 5 + √x) = 1
\Rightarrow√x + 5 + √x = 5 \Rightarrow√x + 5 = 5 - √x
Squaring both sides
\Rightarrowx + 5 = 25 – 10 √x + √x \Rightarrow10√x = 20
\Rightarrow√x = 2 \Rightarrowx = 4

Thanks
Deepak Patra
askIITians Faculty